OpenStudy (anonymous):

Because of friction and air resistance, each swing of a pendulum is a little shorter than the previous one. The lengths of the swings form a geometric sequence. Suppose the first swing of the pendulum has an arc length of 100cm and a return swing of 99cm. a.) On which swing will the length first have a length less than 50 cm? b.) Find the total distance traveled by the pendulum until it comes to rest.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I know the answers but I want to see how they got them.

OpenStudy (anonymous):

a: 70th swing b: 10,000 cm

jimthompson5910 (jim_thompson5910):

you would use the formula an = 100*(0.99)^(n-1) plug in an = 50 to get 50 = 100*(0.99)^(n-1) then solve for n

jimthompson5910 (jim_thompson5910):

this is all for part a

jimthompson5910 (jim_thompson5910):

for part b), you use the formula S = a/(1-r)

OpenStudy (anonymous):

can you show me step by step so ikno i wrote it right? pleasee!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

50 = 100*(0.99)^(n-1) 50/100 = (0.99)^(n-1) 0.5 = 0.99^(n-1) log(0.5) = log(0.99^(n-1)) log(0.5) = (n-1)log(0.99) log(0.5)/log(0.99) = n-1 I'll let you finish

OpenStudy (anonymous):

Did you ever get the rest of this? @jim_thompson5910 Could you finish this so I can check my answer?